A Biotech company produces enzymes at a rate that can be described by the Poisson process.



Their lipase enzyme from a genetically modified bacteria produces glycerol at an average rate of



3 mg/L per minute.



4.1 Find the probability that the enzyme will produce exactly 5 mg/L in a given minute. (2)



4.2 What is the likelihood that the enzyme will produce four or more mg/L in a given minute? (7)



4.3 What chance is there that no product will be produced in a given minute?

1
Expert's answer
2022-04-05T15:58:35-0400

"P(k)=\frac{\lambda^ke^{-\lambda}}{k!}"

"\lambda =3"

1. k=5 "P(5)=\frac{3^5e^{-3}}{5!}=0.1"

2. "P(X \ge 4)=1-P(3)-P(2)-P(1)-P(0)"

"P(3)=\frac{3^3e^{-3}}{3!}=0.224"

"P(2)=\frac{3^2e^{-3}}{2!}=0.224"

"P(1)=\frac{3^1e^{-3}}{1!}=0.149"

"P(0)=\frac{3^0e^{-3}}{0!}=0.05"

"P(X \ge 4)=1-0.224-0.224-0.149-0.05=0.353"

3."P(0)=0.05"